CAN YOU PLEASE HLP ME FIND THE ASYMPTOTES OF Y=(X+1)/(X-5)
Answer by stanbon(75887)

FIND THE ASYMPTOTES OF Y=(X+1)/(X-5)
-------------
Vertical Asymptote:
Condition: denominator = 0
x-5=0
x=5
There is a vertical asymptote at x=5
--------------------
Horizontal Aymptote:
The highest power of x is x^1
The coefficients of x in the numerator and and denominator are 1
There is a horizontal asymptote at y=1/1 or y=1
Cheers,
Stan H.